# Origin of the mm-submm loss in deposited dielectrics

### v1

@@ I Introduction
We measured the loss of hydrogenated amorphous SiC (a-SiC:H) films in the 0.27–100 THz range, where in the 270–600 GHz range we used superconducting microstrip resonators, and above 3 THz we used Fourier-transform spectroscopy (FTS). The agreement between the loss data and a Maxwell-Helmholtz-Drude dispersion model suggests that phonon modes above 10 THz dominate the loss in the a-SiC:H above 200 GHz.
@@ IV.3 Results
Figure 7: a) The points show the transmission through the a-SiC:H samples measured using FTS in the range of 3–100 THz, for S1–3. The solid grey curves show the best-fit results of TMM models where the $\hat{\varepsilon}_{\mathrm{r}}$ is parameterized using the Maxwell-Helmholtz-Drude model (Eq. (3)). The plotted data of S3 was divided by the measured transmission of the bare c-Si substrate to clearly show the absorption features of the a-SiC:H. b) The real ($\varepsilon_{\mathrm{r}}^{{}^{\prime}}$) and imaginary ($\varepsilon_{\mathrm{r}}^{{}^{\prime\prime}}$) parts of the complex permittivity $\hat{\varepsilon}_{\mathrm{r}}$ resulting from the fitting to the transmission data in Fig. 7a. The thin curves show the fitting results in the range of 3-100 THz, and the thick curves show the ranges where measurement data was available. The annotations identify the modes that are present in the absorption spectrum, which are detailed in Appendix A.
Figure 8: Loss tangent $\tan\delta$ versus frequency of the a-SiC:H. The solid grey curve represents the $\tan\delta$ which we obtained by fitting the Maxwell-Helmholtz-Drude model to the combined FTS-measured transmission data (Fig. 7a) and FP-measured loss data (Fig. 5), as described in Section V. The purple curve represents the $\tan\delta$ which we obtained by fitting the Maxwell-Helmholtz-Drude model to only the FTS data (Fig. 7a). The blue points represent the FP-measured loss data from this work (Fig. 5). The green points represent the a-SiC:H loss data reported in Ref. [8]. The intersection of the solid grey curve and the horizontal dashed line illustrates the crossover between the low-power cryogenic TLS loss and phonon mode absorption.
We determined the 3-100 THz loss by fitting transfer matrix method (TMM) [25] models to the FTS transmission data of S1–3. The transmission data and the best-fit transmission curves are presented in Fig. 7a. Here the complex permittivity $\hat{\varepsilon}_{\mathrm{r}}\equiv\varepsilon_{\mathrm{r}}^{{}^{\prime}}+i\varepsilon_{\mathrm{r}}^{{}^{\prime\prime}}$ (and therefore the loss) of the a-SiC:H was a model parameter which we obtained from the fitting. We combined the FTS transmission data of S1–3 in a single fitting procedure, where we normalized the fitting residuals by the number of data points in each data set. In the case of S3, we measured the bare c-Si substrate prior to the deposition of the a-SiC:H and we included the substrate properties in the TMM model. We note that the periodicity in the transmission of S2 (Fig. 7a ) is dictated by the inter-membrane distance.
In the fitting of the TMM models to the transmission data we parameterized the $\hat{\varepsilon}_{\mathrm{r}}$ of the a-SiC:H using the Maxwell-Helmholtz-Drude dispersion model [26, 27]:
@@ V Model of dielectric loss from microwave to mid-infrared
Finally, we examined whether the loss measured with the FPs at 270–600 GHz (Fig. 5) and the loss measured by FTS in the range of 3–100 THz (fig. 7b) can be consistently explained with a common Maxwell-Helmholtz-Drude phonon-loss model. We searched for a set of parameters in Eq. (3) using a numerical optimization algorithm that simultaneously minimizes the differences between (1) the calculated $\tan{\delta}$ and the FP-measured $\tan{\delta}$, and (2) the calculated transmission and the FTS-measured transmission. In the fitting we included the FP-measured $\tan{\delta}$ values from Ref. [8] (Fig. 5). The 520 GHz and 600 GHz points in Fig. 5 had to be excluded because they could not be reproduced by our current model. The resulting best-fit parameter values are listed in Table 2 in Appendix A. In Fig. 8 we co-plotted the $\tan\delta$ resulting from the combined fitting to the FTS and FP measurements (grey curve), and the $\tan\delta$ corresponding to Fig. 7b resulting from FTS measurements alone (purple curves). Except for the points at 520 and at 600 GHz (which will be discussed in the next section), the modeled frequency dependence of $\tan\delta$ agrees well with both the FP and FTS measurements over the entire frequency range that was covered.
@@ VI Discussion on the origin of the loss in deposited dielectrics
Our results show that the absorption tail of phonon modes located at frequencies above 10 THz can explain the measured losses in the range of 270–455 GHz. Above a frequency of roughly 200 GHz the losses are anticipated to be dominated by phonon modes, and below a frequency of roughly 200 GHz and at cryogenic temperatures and low power the loss is expected to be dominated by TLSs. At 520 GHz and at 600 GHz the on-chip loss data is not reproducible by the phonon-loss model with the modes that have been reported in literature. For an overview of the phonon absorption modes in amorphous and crystalline SiC we refer to Refs. [29, 30]. The excess loss at 520 GHz and at 600 GHz is not expected to originate from the bulk of the superconductor since the NbTiN has an energy gap of roughly 1.1 THz. However, it is possible that the excess loss is caused by a small fraction of the superconductor which behaves differently than the bulk of the superconductor. More experiments are needed to clarify this discrepancy between the model and our data at 520 GHz and at 600 GHz.
Our results pave the way for a thorough understanding of the dielectric loss in deposited dielectrics from microwave to near-infrared wavelengths. We emphasize that strong phonon absorption modes in the 10–30 THz range are present not only in a-SiC:H, but have also been reported for other commonly used deposited dielectrics such as a-Si:H [31, 32], $\mathrm{SiN}_{x}$ [26] and $\mathrm{SiO}_{x}$ [27]. Furthermore, the cryogenic and low-power TLS loss tangent $\tan\delta_{\mathrm{TLS}}$ values reported at microwave frequencies of a-Si:H ($\mathrm{\sim}10^{-5}$) [10, 33, 9], $\mathrm{SiN}_{x}$ ($\mathrm{\sim}10^{-4}$) [10] and $\mathrm{SiO}_{x}$ ($\mathrm{\sim}10^{-4}$$10^{-3}$) [10, 34] lead us to anticipate a similar crossover from TLS dominated loss to phonon mode absorption dominated loss for these materials.
@@ VII Conclusion
We demonstrated that the dielectric loss of a-SiC:H above 200 GHz is dominated by the absorption tail of phonon modes, which are located at frequencies above 10 THz. Our results pave the way for a thorough understanding of the dielectric loss in deposited dielectrics from microwave to near infrared wavelengths.
@@ Appendix A FTS fitting parameters
Parameters corresponding to Fig. 7
$j$ Absorption band $\omega_{T,j}/2\pi$ (THz) $\Delta\varepsilon_{j}^{{}^{\prime}}$ $\Gamma_{j}$(THz) $\alpha_{j}$
@@
Table 1: Best-fit values of the fitting parameters corresponding to the results in Fig. 7b, where Eq. 3 was fitted to only the FTS-measured transmission data (Fig. 7a). The $\varepsilon_{\infty}$ resulting from the fit equals 6.8. The values shown in bold were constrained to the listed values to reduce the number of free fitting parameters.
Parameters corresponding to Fig. 8
$j$ Absorption band $\omega_{T,j}/2\pi$ (THz) $\Delta\varepsilon_{j}^{{}^{\prime}}$ $\Gamma_{j}$(THz) $\alpha_{j}$
@@
Table 2: Best-fit values of the fitting parameters corresponding to the solid grey curve in Fig. 8, where Eq. 3 was fitted to the combined FTS-measured transmission data (Fig. 7a) and the FP-measured loss data (Fig. 5). The $\varepsilon_{\infty}$ resulting from the fit equals 6.8. The values shown in bold were constrained to the listed values to reduce the number of free fitting parameters.

### v2

# Vibrational modes as the origin of dielectric loss at 0.27–100 THz in a-SiC:H

@@ I Introduction
We measured the loss of hydrogenated amorphous SiC (a-SiC:H) films in the 0.27–100 THz range, where in the 270–600 GHz range we used superconducting microstrip resonators, and above 3 THz we used Fourier-transform spectroscopy (FTS). The agreement between the loss data and an MHD dispersion model suggests that vibrational modes above 10 THz dominate the loss in the a-SiC:H above 200 GHz.
@@ IV.3 Results
Figure 7: a) The points show the transmission through the a-SiC:H samples measured using FTS in the range of 3–100 THz, for S1–3. The solid grey curves show the best-fit results of TMM models where the $\hat{\varepsilon}_{\mathrm{r}}$ is parameterized using the MHD dispersion model (Eq. (3)). The plotted data of S3 was divided by the measured transmission of the bare c-Si substrate to clearly show the absorption features of the a-SiC:H. b) The real ($\varepsilon_{\mathrm{r}}^{{}^{\prime}}$) and imaginary ($\varepsilon_{\mathrm{r}}^{{}^{\prime\prime}}$) parts of the complex permittivity $\hat{\varepsilon}_{\mathrm{r}}$ resulting from the fitting to the transmission data in Fig. 7a. The thin curves show the fitting results in the range of 3-100 THz, and the thick curves show the ranges where measurement data was available. The annotations identify the modes that are present in the absorption spectrum, which are detailed in Appendix A.
Figure 8: Loss tangent $\tan\delta$ versus frequency of the a-SiC:H. The solid grey curve represents the $\tan\delta$ which we obtained by fitting the MHD dispersion model to the combined FTS-measured transmission data (Fig. 7a) and FP-measured loss data (Fig. 5), as described in Section V. The purple curve represents the $\tan\delta$ which we obtained by fitting the MHD dispersion model (Eq. (3)) to only the FTS data (Fig. 7a). The blue points represent the FP-measured loss data from this work (Fig. 5). The green points represent the a-SiC:H loss data reported in Ref. [8]. The intersection of the solid grey curve and the horizontal dashed line illustrates the crossover between the low-power cryogenic TLS loss and the loss due to vibrational modes.
We determined the 3–100 THz loss by fitting transfer matrix method (TMM) [25] models to the FTS transmission data of S1–3. The transmission data and the best-fit transmission curves are presented in Fig. 7a. Here the complex permittivity $\hat{\varepsilon}_{\mathrm{r}}\equiv\varepsilon_{\mathrm{r}}^{{}^{\prime}}+i\varepsilon_{\mathrm{r}}^{{}^{\prime\prime}}$ (and therefore the loss) of the a-SiC:H was a model parameter which we obtained from the fitting. We combined the FTS transmission data of S1–3 in a single fitting procedure, where we normalized the fitting residuals by the number of data points in each data set. In the case of S3, we measured the bare c-Si substrate prior to the deposition of the a-SiC:H and we included the substrate properties in the TMM model. We note that the periodicity in the transmission of S2 (Fig. 7a ) is dictated by the inter-membrane distance.
In the fitting of the TMM models to the transmission data we parameterized the $\hat{\varepsilon}_{\mathrm{r}}$ of the a-SiC:H using the Maxwell-Helmholtz-Drude (MHD) dispersion model [26, 27]:
@@ V Model of dielectric loss from microwave to mid-infrared
Finally, we examined whether the loss measured with the FPs at 270–600 GHz (Fig. 5) and the loss measured by FTS in the range of 3–100 THz (fig. 7b) can be consistently explained with a common MHD dispersion model. We searched for a set of parameters in Eq. (3) using a numerical optimization algorithm that simultaneously minimizes the differences between (1) the calculated $\tan{\delta}$ and the FP-measured $\tan{\delta}$, and (2) the calculated transmission and the FTS-measured transmission. In the fitting we included the FP-measured $\tan{\delta}$ values from Ref. [8] (Fig. 5). The 520 GHz and 600 GHz points in Fig. 5 had to be excluded because they could not be reproduced by our current model. The resulting best-fit parameter values are listed in Table 2 in Appendix A. In Fig. 8 we co-plotted the $\tan\delta$ resulting from the combined fitting to the FTS and FP measurements (grey curve), and the $\tan\delta$ corresponding to Fig. 7b resulting from FTS measurements alone (purple curve). Except for the points at 520 and at 600 GHz (which will be discussed in the next section), the modeled frequency dependence of $\tan\delta$ agrees well with both the FP and FTS measurements over the entire frequency range that was covered.
@@ VI Discussion on the origin of the loss in deposited dielectrics
Our results show that the absorption tail of vibrational modes located at frequencies above 10 THz can explain the measured losses in the range of 270–455 GHz. Above a frequency of roughly 200 GHz the losses are anticipated to be dominated by vibrational modes, and below a frequency of roughly 200 GHz and at cryogenic temperatures and low power the loss is expected to be dominated by TLSs. At 520 GHz and at 600 GHz the on-chip loss data is not reproducible by the MHD dispersion model with the modes that have been reported in literature. For an overview of the modes in amorphous and crystalline SiC we refer to Refs. [29, 30]. The excess loss at 520 GHz and at 600 GHz is not expected to originate from the bulk of the superconductor since the NbTiN has an energy gap of roughly 1.1 THz. However, it is possible that the excess loss is caused by a small fraction of the superconductor which behaves differently than the bulk of the superconductor. More experiments are needed to clarify this discrepancy between the model and our data at 520 GHz and at 600 GHz.
Our results pave the way for a thorough understanding of the dielectric loss in deposited dielectrics from microwave to near-infrared wavelengths. We emphasize that strong vibrational modes in the 10–30 THz range are present not only in a-SiC:H, but have also been reported for other commonly used deposited dielectrics such as a-Si:H [31, 32], $\mathrm{SiN}_{x}$ [26] and $\mathrm{SiO}_{x}$ [27]. Furthermore, the cryogenic and low-power TLS loss tangent $\tan\delta_{\mathrm{TLS}}$ values reported at microwave frequencies of a-Si:H ($\mathrm{\sim}10^{-5}$) [10, 33, 9], $\mathrm{SiN}_{x}$ ($\mathrm{\sim}10^{-4}$) [10] and $\mathrm{SiO}_{x}$ ($\mathrm{\sim}10^{-4}$$10^{-3}$) [10, 34] lead us to anticipate a similar crossover for these materials, from loss dominated by TLSs to loss dominated by vibrational modes.
@@ VII Conclusion
We demonstrated that the dielectric loss of a-SiC:H above 200 GHz is dominated by the absorption tail of vibrational modes which are located at frequencies above 10 THz. Our results pave the way for a thorough understanding of the dielectric loss in deposited dielectrics from microwave to near infrared wavelengths.
@@ Appendix A FTS fitting parameters
$j$ Vibrational mode $\omega_{T,j}/2\pi$ (THz) $\Delta\varepsilon_{j}^{{}^{\prime}}$ $\Gamma_{j}$(THz) $\alpha_{j}$
@@
Table 1: Best-fit values corresponding to the results in Fig. 7b, of the MHD dispersion model paramaters (Eq. 3). The MHD dispersion model was fitted to only the FTS-measured transmission data (Fig. 7a). The $\varepsilon_{\infty}$ resulting from the fit equals 6.8. The values shown in bold were constrained to the listed values to reduce the number of free fitting parameters.
$j$ Vibrational mode $\omega_{T,j}/2\pi$ (THz) $\Delta\varepsilon_{j}^{{}^{\prime}}$ $\Gamma_{j}$(THz) $\alpha_{j}$
@@
Table 2: Best-fit values corresponding to the solid grey curve in Fig. 8, of the MHD dispersion model parameters (Eq. 3). The MHD dispersion model was fitted to the combined FTS-measured transmission data (Fig. 7a) and the FP-measured loss data (Fig. 5). The $\varepsilon_{\infty}$ resulting from the fit equals 6.8. The values shown in bold were constrained to the listed values to reduce the number of free fitting parameters.